### v1

# Train-before-Test Harmonizes Language Model Rankings

Abstract
Existing language model benchmarks provide contradictory model rankings, even for benchmarks that aim to capture similar skills. This dilemma of conflicting rankings hampers model selection, clouds model comparisons, and adds confusion to a growing ecosystem of competing models. Recent work attributed ranking disagreement to the phenomenon of training on the test task: As released, different models exhibit a different level of preparation for any given test task. A candidate solution to the problem is train-before-test: Give each model the same benchmark-specific finetuning before evaluation. Our primary contribution is a broad empirical evaluation of train-before-test across 24 benchmarks and 61 models. We show that train-before-test significantly improves ranking agreement consistently across all benchmarks. Whereas rankings have little external validity to start with, they enjoy a significant degree of external validity when applying train-before-test: Model rankings transfer gracefully from one benchmark to the other. Even within the same model family, train-before-test reduces strong ranking disagreement to near-perfect agreement. In addition, train-before-test reduces the model-score matrix to essentially rank one, revealing new insights into the latent factors of benchmark performance. Our work supports the recommendation to make train-before-test a default component of LLM benchmarking [^2] Code is available at https://github.com/socialfoundations/lm-harmony..
1 Introduction
A single large language model can reliably solve many tasks. This impressive capability, however, has led to a crisis in benchmarking. A rapidly growing flurry of benchmarks tries to capture the numerous different downstream tasks that we might use a language model for. Unfortunately, different benchmarks generally give contradictory model rankings, making it difficult to compare and choose models.
Traditionally, benchmarking compared models trained for the same task on the same training data. In contrast, most large language models train on proprietary data mixes that differ from one model to the other. Recent work showed that this leads to the problem of training on the test task: the extent to which a model has encountered data similar to the test task during training confounds model comparisons, rankings, and scaling laws kaplan2020scaling . Put simply, an otherwise worse model may have simply prepared better for a specific task.
A candidate solution to the problem is to fine-tune each model on the same task-specific training data before evaluation. We call this form of evaluation train-before-test. We draw a distinction between direct evaluation and train-before-test. In direct evaluation, we evaluate the model as it ships out of the box. Under train-before-test, we fine-tune the model before evaluation. The point of fine-tuning is not to make the model better, but to level the playing field for evaluation. The hope is that train-before-test might level out task-specific model differences so as to reveal more robust model comparisons. Our primary contribution is a broad evaluation of train-before-test, which confirms that this is indeed the case.
1.1 Our Contributions
Figure 1: Rankings of 61 language models on two question-answering benchmarks: Natural Questions Open and ARC Challenge. Left: Direct evaluation leads to inconsistent rankings. Although both benchmarks test for question-answering ability, the resulting model rankings show substantial disagreement. Right: Train-before-test aligns model rankings. Note: For each of the two plots, we greedily align model rankings as much as possible without violating confidence intervals, thus revealing only those ranking changes that are statistically significant. See Appendix C.1 for more details.
In more detail, we make the following key contributions.
Direct evaluation leads to ranking disagreement even between related tasks.
We demonstrate that the prevalent direct evaluation scheme results in strong disagreement between model ranking across various benchmarks. We show that this strong ranking disagreement persists even when restricting to benchmarks that aim to capture similar tasks. Moreover, rankings still strongly disagree when evaluating models from the same family. The situation presents a serious conundrum for model selection: Under direct evaluation, benchmarks fail to give reliable and actionable insights about which model to choose from a set of available alternatives.
Train-before-test leads to ranking agreement.
We broadly evaluate train-before-test on 24 benchmark datasets and 61 large language models. Train-before-test will fine-tune each model on the same task-relevant data before evaluation. Our results demonstrate that train-before-test drastically increases model ranking agreement across the board: Ranking agreement between benchmarks, measured by Kendall’s tau, improves for $95\%$ of the $\binom{24}{2}=276$ benchmark pairs. The average value of Kendall’s $\tau$ improves from $0.51$ to $0.74$. Figure 1 illustrates the result for one typical pair of benchmarks.
Train-before-test aligns perplexity rankings with downstream tasks.
Perplexity benchmarks used to be popular, but fell out of fashion because of the apparent disconnect between perplexity and downstream task performance (weiemergent, ; ganguli2022predictability, ; liu2023same, ). We indeed validate this disconnect when comparing model families under direct evaluation. In contrast, we demonstrate that under train-before-test evaluation, perplexity rankings align well with downstream task rankings. This creates an appealing consistency between model training objectives and downstream performance.
Train-before-test sheds light on the latent factors of benchmark performance.
Consider the large benchmark-model matrix, where each entry $(i,j)$ corresponds to the performance of model $j$ on a benchmark $i$. Several works have considered this matrix for different reasons and found that it is approximately low rank (Ruan2024ObservationalSL, ; Owen2024HowPI, ; Burnell2023RevealingTS, ), but not quite. The first singular value is dominant and correlates with pre-training compute. However, the other components aren’t negligible and it wasn’t clear how to interpret them.
We show that train-before-test clarifies this state of affairs. After train-before-test, the benchmark-model matrix is essentially rank $1$. The first principal component accounts for $85\%$ of the explained variance across all models, and for $93\%$ of the variance for a single model family. This leads us to speculate that the lower components of the direct evaluation matrix correspond to how much a model was trained on a given kind of task. This would be a compelling reason why the lower singular values shrink under train-before-test.
2 Related Work
Benchmarking has played a central role in the advancement of machine learning (liberman2010obituary, ; hardt2022patterns, ). While absolute model performance is often fragile to even seemingly minor changes in evaluation data (candela2009dataset, ; torralba2011unbiased, ; albadawy2018deep, ; taori2020measuring, ; Tsipras2020FromIT, ; miller2020effect, ), relative model performance—that is, model rankings—tends to transfer surprisingly well across classical benchmarks (Yadav2019ColdCT, ; recht2019imagenetclassifiersgeneralizeimagenet, ; miller2020effect, ). For instance, prior work (kornblith2019better, ; barbu2019objectnet, ) has shown that model rankings on ImageNet (deng2009imagenet, ) also transfer to other image classification and object recognition benchmarks. Moreover, Salaudeen and Hardt (2024, (salaudeen2024imagenot, )) demonstrated that ImageNet rankings remain robust even under major dataset variations (salaudeen2024imagenot, ). This transferability of model rankings is highly desirable, as it indicates that progress on specific benchmarks reliably reflects broader scientific advancements (liao2021we, ; hardt2025emerging, ).
However, the emergence of foundation models has dramatically transformed the benchmarking landscape compared to the ImageNet era (liang2022holistic, ; srivastava2022beyond, ; Weidinger2025TowardAE, ). With huge training costs and much improved capabilities (qwen, ; gemma, ; llama, ; Ramesh2021ZeroShotTG, ; Anil2023GeminiAF, ; openai2023GPT4, ), practitioners now lean towards directly evaluating LLMs across a wide range of different benchmarks, in the hope of obtaining a more comprehensive assessment of their capabilities (liang2022holistic, ; bbh, ; mmlu, ; open-llm-leaderboard, ; open-llm-leaderboard-v2, ). This shift introduces new challenges, as model rankings across different tasks may vary significantly (liang2022holistic, ; huan2025does, ; lourie2025scaling, ). Zhang and Hardt (2024, (zhang2024inherent, )) draw an analogy between multi-task benchmarks and voting systems (arrow2012social, ), revealing that a multi-benchmarking approach with diverse rankings inherently lacks robustness to minor changes and thus cannot provide a stable unified ranking.
This lack of unified ranking is sometimes seen as a desirable feature within the community (liang2022holistic, ). Some argue that variability reflects the multifaceted strengths and weaknesses of LLMs, suggesting that users should select the best model tailored to their specific needs (ghosh2024onebench, ; Zhang2024TaskMA, ; Shnitzer2023LargeLM, ). For example, a user who focuses on mathematical tasks could prioritize the math benchmark to choose the optimum model. However, there are two significant concerns regarding this approach: First, the user-driven selection strategy poses challenges for model developers. Given the resource-intensive nature of LLM development (guo2025deepseek, ), it is impractical to release a different model for every potential use case. Moreover, developers typically aim to create a general-purpose model (qwen, ; llama, ; gemma, ); however, such a desideratum is often difficult to reliably measure due to the inconsistent rankings observed across benchmarks. Second, we demonstrate in this paper that benchmarks within the same task category can still exhibit substantial discrepancies in model rankings.
One potential reason for the observed inconsistencies in model rankings is that models vary substantially in their training data (Gadre2023DataCompIS, ; albalaksurvey, ). In particular, Dominguez et al. (2024, (dominguez2024training, )) show that models vary in their degree of preparedness for popular benchmarks. Building on this idea, we introduce the notion of train-before-test, wherein we fine-tune each model on the corresponding training set to get every model well-prepared. We then investigate how this approach improves the consistency of rankings across benchmarks and discuss its implications for future benchmarking practices.
3 Experiments
In this section, we examine the cross-benchmark ranking agreement of 61 language models across 24 benchmarks. We find that ranking agreement tends to be low, with an average Kendall’s $\tau$ of 0.51. We then examine benchmark agreement under a different benchmarking methodology, which we refer to as train-before-test. Specifically, we fine-tune on a benchmark’s train set prior to evaluating on said benchmark. Compared to direct evaluation, train-before-test improves cross-benchmark ranking agreement on 95% of the benchmark pairs considered. The improvements in ranking agreement are typically large, with the average Kendall’s $\tau$ increasing to 0.74.
We additionally find that train-before-test greatly improves the agreement between perplexity rankings and downstream benchmarks. This result holds consistently across three general domain corpus, newly collected from Wiki, arXiv, and StackExchange. The average Kendall’s $\tau$ improves from 0.47 to 0.73 with train-before-test, leading to much better consistency between the training objective and downstream benchmark performance.
Finally, we discuss the implications of increased cross-benchmark agreement. In doing so, we perform Principal Component Analysis (PCA) over the model score matrix comprising all benchmark scores. We then analyze its principal components both under direct model evaluation and train-before-test. We find that train-before-test greatly increases the share of variance explained by the first principal component (PC1), from 69% for direct evaluation to 85% for train-before-test. In both cases, PC1 aligns well with model pre-training compute. By further controlling for model family and only considering Qwen models, we show that the explained variance ratio of PC1 further increases to 93%, making the model-score matrix essentially rank one.
Table 1: We categorize benchmarks into language understanding (LU), commonsense reasoning (CR), question answering (QA), physics/biology/chemistry (PBC), math (Math), and medicine (Med).
Category Benchmarks
LU
MNLI (mnli, ), QNLI (qnli, ), RTE (rte, ; rte2, ; rte3, ), CoLA (cola, ), SST-2 (sst2, ), MRPC (mrpc, ), QQP, WiC (wic, ), ANLI (anli, )
CR
Winogrande (winograd, ), CommonsenseQA (commonsenseqa, ), Hellaswag (hellaswag, ), Social-IQA (socialiqa, )
QA
OpenBookQA (openbookqa, ), NQ-Open (nq_open, ), BoolQ (boolq, ), ARC-Easy, ARC-Challenge (allenai:arc, )
PBC SciQ (sciq, ), PIQA (piqa, )
Math MathQA (mathqa, ), GSM8K (gsm8k, )
Med MedMCQA (medmcqa, ), HeadQA (headqa, )
Figure 2: Mean ranking agreement between each benchmark and all others. We calculate Kendall’s $\tau$ between each benchmark and every other benchmark, and then average the results. Compared to direct evaluation, train-before-test consistently improves ranking agreement, often by a large margin. A detailed comparison of Kendall’s $\tau$ values for every benchmark pair is provided in Appendix B.1. On average, the overall average Kendall’s $\tau$ is 0.51 for direct evaluation and 0.74 for train-before-test.
(a) Direct evaluation.
(b) Train-before-test.
Figure 3: Cross-category ranking agreement for direct evaluation (left) and train-before-test (right). We categorize benchmarks into language understanding (LU), commonsense reasoning (CR), question answering (QA), physics/biology/chemistry (PBC), math (Math), and medicine (Med), see Table 1. Kendall’s $\tau$ is averaged across all pairs of benchmarks that belong to two given categories. The diagonal entries represent intra-category agreement and the other entries represent inter-category agreement. Train-before-test improves both intra- and inter-category ranking agreement in all instances.
3.1 Experiment Setting
Benchmark selection.
We begin our study with the lm-eval-harness package (eval-harness, ), which offers a comprehensive suite of language model benchmarks. To accommodate the train-before-test methodology which requires a dedicated training set for fine-tuning, we first identify benchmarks that provide at least 1,000 training examples. This yields a total of 37 benchmarks, which we broadly categorize into 28 likelihood-based and 9 generation-based benchmarks.
Generation-based benchmarks are often computationally intensive to evaluate, as base models typically generate text until reaching their maximum sequence length. We select only NQ-Open and GSM8K from the generation-based benchmarks, excluding the remaining seven due to computational constraints. Among the likelihood-based benchmarks, we further exclude six due to observed anomalies during fine-tuning, such as a lack of performance improvement in over 20% of models. See Appendix A.1 for additional details.
Our final selection consists of 24 benchmarks covering diverse domains and task types. These benchmarks are primarily multiple-choice question answering benchmarks, with accuracy as the task metric. We categorize all benchmarks by their descriptions, see Table 1.
If a benchmark does not come with a validation split, we randomly allocate 20% of the training data as the validation set. To save computational resources, we cap the number of training data at 50,000, validation data at 1,000, and testing data at 10,000.
Model selection.
We consider 61 language models across six model families: Llama (llama, ), Qwen (qwen, ), Gemma (gemma, ), Pythia (pythia, ), GPT-2 (gpt2, ) and Yi (yi, ). Due to computational constraints, we select models with fewer than 14B parameters. See Table 2 for the full list. We include both base and instruction-tuned models, and use the suffix -IT to denote instruction-tuned models.
Evaluation setup.
We evaluate the 61 models across all 24 benchmarks using both direct evaluation and train-before-test evaluation. We use the lm-eval-harness library for evaluation. We evaluate models zero-shot (Brown2020LanguageMA, ). For direct evaluation, we simply evaluate the model as it is. For train-before-test, we fine-tune models for five epochs, and select the best performing checkpoint based on performance on a separate validation set. We therefore fine-tune a total of $61$ models on $24$ benchmarks, totaling $61\times 24=1,464$ fine-tuned models. We use parameter-efficient fine-tuning (PEFT) (Hu2021LoRALA, ; peft, ), with fixed hyperparameters. See more details in Appendix A.3. Each fine-tuned model is then evaluated on the corresponding benchmark’s test set. For each benchmark, we rank models according to their performance. We then measure the ranking correlation across pairs of benchmarks using Kendall’s $\tau$ (kendall1938new, ).
3.2 Downstream Ranking Agreement
As depicted in Figure 2, direct evaluation shows only modest ranking agreement between the 24 benchmarks, with an average Kendall’s $\tau$ ranking correlation of 0.51. This lack of agreement across benchmarks complicates model assessment and makes it challenging to aggregate results into a meaningful overall ranking. In contrast, the train-before-test methodology leads to a substantial improvement in ranking agreement. Under this approach, 95% of model pairs show higher Kendall’s $\tau$ scores, with the average $\tau$ rising from 0.51 to 0.74. This improvement in ranking consistency is highly notable, as it indicates that models that excel on any given benchmark are also likely to excel in many others, thereby simplifying the model selection process for developers. Notably, benchmarks that appeared to be outliers under direct evaluation, such as MRPC and WiC, demonstrate much greater ranking consistency under train-before-test. For example, the average Kendall’s $tau$ between MRPC and all other benchmarks improves dramatically, from 0.23 to 0.73.
We further split all benchmarks into six categories (e.g., language understanding, math), see Table 1. For each category pair, we report in Figure 3 the intra-category average ranking correlations and inter-category average ranking correlations across all relevant benchmark pairs. Consistent with our previous findings, we observe reasonably poor ranking agreements across benchmark categories. While one might expect high intra-category agreement–after all, tasks within the same category tend to be relatively similar–direct evaluation results in low intra-category agreement in many cases. For example, the intra-category mean Kendall’s $\tau$ is 0.52 for language understanding and 0.56 for math. This further underscores the difficulty of selecting models based on direct evaluation. Even if the goal is to choose a model that excels not across all tasks but within a specific domain, the low intra-category agreement makes this decision challenging.
In contrast, train-before-test markedly boosts both intra- and inter-category consistency. For example, the intra-category mean Kendall’s $\tau$ for language understanding raises from 0.52 to 0.71, as well as from 0.56 to 0.82 for the math category. Moreover, agreement between categories is often nearly as high as agreement within categories. This suggests that models performing well in one domain tend to also perform well across other domains.
3.3 Perplexity Agreement
We now study the agreement between benchmark rankings and rankings based on perplexity on general domain corpora. To do so, we collect three corpora from Wikipedia, StackExchange, and arXiv, retaining only contents from 2025. More specifically, we collect 3,366 documents for Wiki, 6,001 for StackExchange and 44,384 documents for arXiv. These datasets are split into training, validation, and testing sets, in an 8:1:1 ratio.
Figure 4: Ranking agreement between perplexity rankings and downstream benchmark rankings under direct evaluation (top) and train-before-test (bottom). Perplexity rankings are consistent with each other under both evaluation schemes, with an average Kendall’s $\tau$ of 0.76 and 0.78, respectively. However, for direct evaluation, agreement between perplexity rankings and downstream rankings is low, with an average Kendall’s $\tau$ of just 0.47. Fortunately, train-before-test results in much higher agreement between perplexity and downstream evaluations, increasing average Kendall’s $\tau$ to 0.73.
Figure 5: Ranking agreement between perplexity rankings before fine-tuning and downstream benchmark rankings after fine-tuning, for base models (top) and instruction-tuned models (bottom). For base models, perplexity rankings before fine-tuning are consistent with downstream performance rankings after fine-tuning, with an average of Kendall’s $\tau$ 0.77. The same does not hold for instruction-tuned models, where the average Kendall’s $\tau$ between perplexity rankings and fine-tuning rankings is 0.5.
We measure perplexity in bits per byte, using the lm-eval-harness library. We then construct models rankings based on the perplexity evaluations, and contrast the rankings with those of the downstream benchmarks considered in earlier sections. We exclude the four Gemma models from these results, as lm-eval-harness provides unreliable perplexity measurements for Gemma models due to its rolling window implementation. See Appendix B.2 for more details.
The results are presented in Figure 4. In contrast to downstream tasks, perplexity rankings demonstrate strong agreement both for direct evaluation and train-before-test. Specifically, the average Kendall’s $\tau$ between the perplexity rankings is 0.76 for direct evaluation and 0.78 for train-before-test. We hypothesize that this reasonably strong agreement arises due to the smooth relationship between perplexity evaluations (brandfonbrener2024loss, ; mayilvahananllms, ).
When comparing ranking agreement between perplexity evaluations and downstream benchmarks, we find that agreement is low under direct evaluation, with a mean Kendall’s $\tau$ of 0.47. This lack of agreement is concerning, as it signals a disconnect between the language modelling pre-training objective and downstream benchmark performance. Fortunately, we find that our train-before-test methodology improves ranking agreement substantially, with the mean Kendall’s $\tau$ ranking correlation between perplexity rankings and benchmark rankings rising to 0.73. This finding is reassuring: a light amount of fine-tuning on task data is sufficient to align the language modeling training objective with downstream performance. Moreover, we find that ranking agreement between perplexity and downstream evaluations is roughly similar to agreement across downstream evaluations. This suggests that, despite perplexity typically not being used for benchmarking purposes, it can be as effective a ranking metric as benchmark evaluations.
Drawing inspiration from prior work (liu2023same, ; xia2023training, ; Gadre2024LanguageMS, ; benchmarkloss, ), we further examine the correlation between model rankings according to average perplexity across the three text corpora and average downstream performance across the 24 benchmarks. Gadre et al. (2024, (Gadre2024LanguageMS, )) show that, when models are trained on the same pretraining data, perplexity is well-correlated with aggregate benchmark performance. Our setup differs in that we consider a diverse set of model families, each trained on different pretraining data. Under direct evaluation, we find that the ranking correlation is modest, with a Kendall’s $\tau$ of only 0.56. We hypothesize that this relatively weak agreement is due to differences in pretraining data and instruction tuning, resulting in varying levels of exposure to benchmark tasks during training (dominguez2024training, ). Fortunately, when applying our train-before-test methodology, the ranking correlation between average perplexity and average downstream performance improves substantially, with Kendall’s $\tau$ increasing from 0.56 to 0.82.
We additionally examine the agreement between perplexity prior to fine-tuning and downstream task performance after fine-tuning. That is, between direct evaluation perplexity rankings and train-before-test downstream performance rankings. We plot such ranking agreement in Figure 5, dividing models into base models and instruction-tuned models. We find that, for base models, perplexity prior to fine-tuning is a strong indicator of task performance after fine-tuning. This indicates that, for base models, direct evaluation of perplexity is already a reasonably reliable metric for ranking models. However, the same does not hold for instruction-tuned models. We find that instruction-tuning renders perplexity rankings unreliable, as ranking agreement is low across the board. This is to be expected: instruction fine-tuning tends to increase both benchmark performance and perplexity on general text corpora, thus clouding the relationship between perplexity and downstream evaluations. Fortunately, as shown earlier, train-before-test restores high ranking agreement between perplexity evaluations and downstream task performance.
Figure 6: Explained variance ratios of the top five principal components of the benchmark score matrix, under direct evaluation (left) and train-before-test (right). Train-before-test substantially increases the amount of variance explained by the first principal component, from 69% to 85%
3.4 Low-Ranked Model Score Matrix
So far, we have shown that the train-before-test methodology significantly improves cross-benchmark ranking agreement. We now examine the implications of this finding by analyzing the resulting matrix of model scores, where each entry $(i,j)$ corresponds to the performance of model $j$ on a benchmark $i$. We use Principal Component Analysis (PCA) to examine the structure of the matrix of model scores.
Figure 6 shows the explained variance ratios of the first five principal components. These results support previous findings that the score matrix is of low rank (Ruan2024ObservationalSL, ). Under direct evaluation, the first five components account for 91% of the total variance. A similar trend is observed for train-before-test scores, where the first five components explain 97% of the variance. Notably, under train-before-test, the first principal component (PC1) captures a significantly larger share of the variance: 85%, compared to 69% for direct evaluation.
Prior works interpret PC1 scores as an indication of general capability, with later principal components denoting domain-specific capabilities not captured by PC1 (e.g., reasoning ability, coding ability) (Ruan2024ObservationalSL, ; Burnell2023RevealingTS, ). Contrary to the popular notion that the performance of LLMs is governed by multiple axes of capability (Ruan2024ObservationalSL, ; Burnell2023RevealingTS, ), our results highlight the dominance of a single principal axis of general capability.
Figure 7: PC1 scores under train-before-test align with the pre-training compute.
It is of no surprise that PC1 positively correlates with pre-training compute, as shown in Figure 7222We only plot models whose number of training tokens is publicly available. See Table 4 for details., which have been identified as crucial to model performances (kaplan2020scaling, ). However, the exact interpretation of PC1 remains open to debate. It could be model scale (kaplan2020scaling, ), general capability (Ruan2024ObservationalSL, ), or another factor entirely. Nonetheless, it is evident that PC1 is strongly linked with model performance–particularly under the train-before-test methodology– and thus might serve as an effective target for guiding further model development. See detailed PC1 scores in Figure 11.
Case study for Qwen models.
Figure 8: Explained variance ratios of the top five principal components of the Qwen score matrix. For train-before-test, the explained variance ratio of PC1 increases to 93%, making the Qwen score matrix essentially rank one.
We repeat our PCA analysis on the score matrix containing only Qwen models, depicted in Figure 8. Remarkably, we find that PC1 for train-before-test explains 93% of the variance, roughly as much as the variance explained by the top four principal components under direct evaluation. That is, whereas for direct evaluation the score matrix is low-rank; train-before-test renders the score matrix essentially rank one.
4 Discussion, limitations, and conclusion
Train-before-test harmonizes model rankings across numerous benchmarks we evaluated. Put differently, it endows model rankings with external validity they lack under direct evaluation. In turn, this greatly increases the utility of benchmarks for reliable model comparisons and model selection. To play devil’s advocate, however, one might ask why do we need ranking consistency if we can just look at the benchmark that best fits a downstream application? First, we show that even benchmarks that aim to capture the same task give contradictory rankings. Second, there will never be a perfect match between a benchmark and the downstream application. We must therefore always rely on some form of external validity, specifically, that the ranking transfers gracefully to other settings. Third, practitioners typically specialize any chosen model to their application. This makes direct evaluation numbers irrelevant, anyway.
Limitations.
Train-before-test requires that we fine-tune models on agreed upon task-specific data prior to evaluation. This certainly increases the cost of evaluation and might be too costly in some cases. On the other hand, it also points to interesting avenues for future work on how to implement train-before-test most efficiently. A second problem is that, unfortunately, many benchmarks no longer come with training data, making it more difficult to apply train-before-test. In light of our findings, we recommend that future benchmarks provide fine-tuning data for the benchmark. A third limitation is that some commercial model providers do not easily allow fine-tuning of their models. We contend that in this case the problem is with the model provider. There is clearly scientific value in creating an ecosystem of models that can be fine-tuned. Train-before-test evaluation creates additional incentives for making models easy to fine-tune.
Conclusion.
On the whole, we believe that train-before-test is a sound recommendation for the evaluation ecosystem that can help address some of its serious problems.
Acknowledgement
We thank Yatong Chen, Florian Dorner, Mina Remeli and Jiduan Wu for helpful discussions and feedback on draft versions of this work.
Appendix A Additional Experiment Setting
A.1 Benchmark Selection
Language model evaluations can be broadly classified into likelihood-based and generation-based. Likelihood-based evaluations test for the likelihood of different completions given some input string; for example, different answer choices given a multiple-choice input question. Since the number of completions is usually small, likelihood-based evaluations are generally compute-efficient. Generation-based evaluations, in contrast, generate some output response given an input query. If responses tend to be long, then generation-based evaluations naturally become compute-intensive. This is particularly true for base models, which are usually not trained for instruction following and therefore continue to generate tokens until hitting their maximum token limit.
Due to computational constraints, we exclude seven generation-based benchmarks, Drop, CoQa, ReCoRD, bAbi, WebQA, TriviaQA and Fld-Default. Nevertheless, we retain two widely used generation-based benchmarks, GSM8K and NQ-Open, in our experiments.
We additionally excluded five benchmarks due to anomalies observed during fine-tuning: MedQA-4Options, LogiQA, Mutual, Mela-EN, and Swag. For these benchmarks, more than 20% of models showed no performance improvement after fine-tuning.
We also excluded Paws-EN, as its corresponding model ranking under direct evaluation was negatively correlated (Kendall’s $\tau$ less than zero) with 23 out of 24 other benchmarks. We attribute this anomaly to the unusual prompting template used by lm-eval-harness.
A.2 Model Selection
See Table 2 for the complete list of models used in our experiments.
Table 2: Models considered, categorized by model family.
Family Model Name Suffix
Llama- [29] 3-8B, 3.1-8B, 3.2-1B, 3.2-3B, 3-8B-IT, 3.1-8B-IT, 3.2-1B-IT, 3.2-3B-IT
Qwen- [56] 1.5-0.5B, 1.5-1.8B, 1.5-4B, 1.5-7B, 1.5-14B, 2-0.5B, 2-1.5B, 2-7B, 2.5-0.5B, 2.5-1.5B, 2.5-3B, 2.5-7B, 2.5-14B, 1.5-0.5B-IT, 1.5-1.8B-IT, 1.5-4B-IT, 1.5-7B-IT, 1.5-14B-IT, 2-0.5B-IT, 2-1.5B-IT, 2-7B-IT, 2.5-0.5B-IT, 2.5-1.5B-IT, 2.5-3B-IT, 2.5-7B-IT, 2.5-14B-IT
Gemma- [71] 2B, 7B, 2-2B, 2-9B, 2B-IT, 7B-IT, 2-2B-IT, 2-9B-IT
GPT2- [57] 124M, 335M, 774M, 1.5B
Pythia- [8] 70M, 160M, 410M, 1B, 1.4B, 2.8B, 6.9B, 12B
Yi- [82] 6B, 9B, 6B-IT, 1.5-6B, 1.5-9B, 1.5-6B-IT, 1.5-9B-IT
A.3 Evaluation Setup
For our train-before-test evaluations, we fine-tune each model for five epochs and select the best-performing checkpoint based on evaluations on a separate validation set. To reduce memory consumption, we employ parameter-efficient fine-tuning (PEFT) [34, 47], allowing us to conduct most experiments on Quadro RTX 6000, Tesla V100-SXM2-32GB and NVIDIA A100-SXM4-80GB GPUs.
We use fixed hyperparameters across all experiments. These hyperparameters were not tuned for any particular model or benchmark. Specifically, we use AdamW with a learning rate of 2e-5 and weight decay of 0.01. For PEFT, we use a LoRA configuration with rank 8, $\alpha=32$, and dropout 0.1.
For models that do not exhibit performance improvements after fine-tuning, we simply report their pre-fine-tuning results. Since all training datasets considered in our work are publicly available, models may have already trained on them during pre-training. In certain cases, this may lead to further fine-tuning not being beneficial.
For instruction-tuned models, we assess performance both with and without applying chat templates and use the best performing one under direct evaluation. In the train-before-test setting, we similarly fine-tune two variants: one where the training data are formatted with chat templates and one without. We similarly select the best performing approach. We observe that not using a chat template often leads to better performance than using a chat template, both for direct evaluation and train-before-test.
Appendix B Additional Experiment Results
B.1 Downstream Ranking Agreement
We plot detailed pairwise ranking correlation agreement between benchmarks in Figure 9 (direct evaluation) and 10 (train-before-test), corresponding to Figure 2 in the main text.
Figure 9: Cross benchmark ranking agreement under direct evaluation. Benchmarks are sorted based on the training dataset size. Kendall’s $\tau$ is calculated for every benchmark pair.
Figure 10: Cross benchmark ranking agreement under train-before-test. Benchmarks are sorted based on the training dataset size. Kendall’s $\tau$ is calculated for every benchmark pair.
B.2 Perplexity Ranking Agreement
Table 3: Bits per byte (BPB) of eight excluded Gemma models compared to Pythia-410M across the three newly collected corpora. The Gemma models exhibit abnormally high BPB values on Wiki and Stack, likely due to the greater average sequence length in these two datasets. Specifically, Arxiv has an average of 163 words per document, compared to 250 for Stack and 1502 for Wiki.
Arxiv Wiki Stack
Gemma-2B 0.766 1.578 1.139
Gemma-2B-IT 0.770 1.524 1.222
Gemma-7B 1.013 4.780 4.053
Gemma-7B-IT 1.053 18.711 20.958
Gemma-2-2B 0.730 1.784 1.340
Gemma-2-2B-IT 0.705 1.191 0.997
Gemma-2-9B 0.709 2.216 1.685
Gemma-2-9B-IT 0.638 1.234 0.978
Pythia-410M 0.791 1.065 0.945
In this work, we collect three corpora from Wikipedia, StackExchange, and arXiv. We only collect documents from 2025. More specifically, we collect 3,366 documents for Wiki, 6,001 for StackExchange and 44,384 documents for arXiv. These datasets are split into training, validation, and testing sets, in an 8:1:1 ratio. For arXiv, we utilize only the paper abstracts, while for StackExchange, we use only the questions. Consequently, the average document length is 163 words for arXiv, 250 words for StackExchange, and 1,502 words for Wikipedia.
We exclude Gemma models from our perplexity agreement experiments, as lm-eval-harness provides unreliable perplexity measurements for Gemma models333 See discussion at https://github.com/huggingface/transformers/issues/29250. . We report the bits per byte (BPB) for the Gemma models in Table 3. While the BPB values for Gemma on arXiv (the dataset with the shortest average sequence length) are mostly reasonable, the performance on StackExchange and Wikipedia is notably worse, even compared to smaller models like Pythia-410M.
This anomaly stems from how lm-eval-harness handles long sequences via a rolling window mechanism. Unlike other models, Gemma requires every input sequence to begin with the BOS token. If this constraint is not met, perplexity degrades significantly. Consequently, when processing long sequences that are chunked into multiple windows, Gemma’s performance degrades.
B.3 PC1 Score under Train-before-Test
We plot the PC1 scores under train-before-test in Figure 11. We also provide the pre-training compute details for models with publicly available training token counts, as shown in Table 4.
Figure 11: PC1 scores under train-before-test.
[…]
Model #Parameters (B) #Tokens (B) #FLOPs (10^18)
Llama-3-8B 8.03 15000.0 722700.00
Llama-3-8B-IT 8.03 15000.0 722700.00
Llama-3.1-8B 8.03 15000.0 722700.00
Llama-3.1-8B-IT 8.03 15000.0 722700.00
Llama-3.2-3B 3.21 9000.0 173340.00
Llama-3.2-3B-IT 3.21 9000.0 173340.00
Qwen1.5-0.5B 0.62 2400.0 8928.00
Qwen1.5-1.8B 1.84 2400.0 26496.00
Qwen1.5-4B 3.95 2400.0 56880.00
Qwen1.5-7B 7.72 4000.0 185280.00
Qwen1.5-14B 14.20 4000.0 340800.00
Qwen1.5-0.5B-IT 0.62 2400.0 8928.00
Qwen1.5-1.8B-IT 1.84 2400.0 26496.00
Qwen1.5-4B-IT 3.95 2400.0 56880.00
Qwen1.5-7B-IT 7.72 4000.0 185280.00
Qwen1.5-14B-IT 14.20 4000.0 340800.00
Gemma-7B 8.54 6000.0 307440.00
Gemma-7B-IT 8.54 6000.0 307440.00
Gemma-2-2B 2.61 2000.0 31320.00
Gemma-2-2B-IT 2.61 2000.0 31320.00
Gemma-2-9B 9.24 8000.0 443520.00
Gemma-2-9B-IT 9.24 8000.0 443520.00
Pythia-70M 0.07 300.0 126.00
Pythia-160M 0.16 300.0 288.00
Pythia-410M 0.41 300.0 738.00
Pythia-1B 1.00 300.0 1800.00
Pythia-1.4B 1.40 300.0 2520.00
Pythia-2.8B 2.80 300.0 5040.00
Pythia-6.9B 6.90 300.0 12420.00
Pythia-12B 12.00 300.0 21600.00
Yi-6B 6.06 3000.0 109080.00
Yi-6B-IT 6.06 3000.0 109080.00
Yi-9B 8.83 3800.0 201324.00
Yi-1.5-6B 6.06 3600.0 130896.00
Yi-1.5-6B-IT 6.06 3600.0 130896.00
Yi-1.5-9B 8.83 3600.0 190728.00
Yi-1.5-9B-IT 8.83 3600.0 190728.00
Appendix C Accounting for Statistical Significance
C.1 Ranking Alignment in Figure 1
We plot the rankings of 61 language models on two question-answering benchmarks: Natural Questions Open and ARC Challenge in Figure 1. We greedily align each ranking as much as possible without violating confidence intervals, thus revealing only those ranking changes that are statistically significant. See Algorithm 3 for more details.
C.2 Downstream Ranking Agreement
We additionally supplement the experiments presented in the main text by modifying the ranking correlation metric to account for statistical significance in the benchmark evaluations. Specifically, we use Kendall’s $\tau$-b [38], which adjusts for ties in rankings. We consider two models tied on a given benchmark if their performance difference is not statistically significant at the 95% confidence level. We assess statistical significance using a t-test based on the standard error of the mean performances.
We reproduce the ranking correlation figures of the main text using the modified Kendall’s $\tau$ which treats non-statistically significant performance differences as ties. See Figure 12 and 13; as well as Figure 14 and Figure 15 for more detailed results. We observe that accounting for statistical significance in models’ performance differences leads to slightly higher ranking correlations, as measured by Kendall’s $\tau$-b. For direct evaluation, average agreement increases from 0.51 to 0.57. For train-before-test, average agreement increases from 0.74 to 0.77. Therefore, train-before-test continues to lead to large improvements in raking agreement (from Kendall’s $\tau$-b 0.57 to 0.77).
Input: Model performance scores and standard errors
Output: Directed graph $G$ representing significant model orderings
Initialize graph $G$ with models as nodes
foreach pair of distinct models $(m_{1},m_{2})$ do
       if $m_{1}$ is significantly better than $m_{2}$ then
             Add directed edge $(m_{1}\rightarrow m_{2})$ to $G$
return $G$
Algorithm 1 build_partial_order(scores, stderrs)
Input: List of models, two directed graphs $G_{1}$, $G_{2}$, and two score series
Output: Two lists representing the parallel ranking order for each task
Initialize vanillaRank1, $\leftarrow$ rankdata(score1), vanillaRank2 $\leftarrow$ rankdata(score2)
Initialize available1 and available2 as models with zero in-degree in $G_{1}$ and $G_{2}$
Initialize empty lists order1, order2
for $i=1$ to number of models do
       Initialize empty list pairs
       foreach $m_{1}$ in available1 do
             foreach $m_{2}$ in available2 do
                   Compute cost for pair $(m_{1},m_{2})$ based on:
                   (1) Placement of $m_{1}$ in order2 and $m_{2}$ in order1
                   (2) Whether $m_{1}=m_{2}$ (prefer matching)
                   (3) Combined vanilla ranks: vanillaRank2[$m_{1}$] + vanillaRank1[$m_{2}$]
                   Append $(cost,m_{1},m_{2})$ to pairs
      Sort pairs by cost (ascending)
       Select $(m_{1},m_{2})$ with minimal cost
       Append $m_{1}$ to order1, $m_{2}$ to order2
       Remove $m_{1}$ from $G_{1}$ and update available1
       Remove $m_{2}$ from $G_{2}$ and update available2
return order1, order2
Algorithm 2 parallel_greedy_rank(models, $G_{1}$, $G_{2}$, score1, score2)
Input: Scores and standard errors for two tasks
Output: Parallel rankings for both tasks
$G_{1}\leftarrow$ build_partial_order(score1, stderr1)
$G_{2}\leftarrow$ build_partial_order(score2, stderr2)
$(order_{1},order_{2})\leftarrow$ parallel_greedy_rank(models, $G_{1}$, $G_{2}$, score1, score2)
rank1[$m$] $=$ position of $m$ in $order_{1}$ (starting from 1)
rank2[$m$] $=$ position of $m$ in $order_{2}$ (starting from 1)
return rank1, rank2
Algorithm 3 rank_models(score1, stderr1, score2, stderr2)
Figure 12: Mean ranking agreement between each benchmark and all others, measured by Kendall’s $tau$-b, with non-statistically significant performance differences being treated as ties. We calculate Kendall’s $\tau$-b between each benchmark and every other one, and then average. Compared to direct evaluation, train-before-test consistently improves ranking agreement–often by a large margin. On average, the overall average Kendall’s $\tau$ is 0.57 for direct evaluation and 0.77 for train-before-test.
(a) Direct evaluation.
(b) Train-before-test.
Figure 13: Cross-category ranking agreement for direct evaluation (left) and train-before-test (right), measured by Kendall’s $tau$-b, with non-statistically significant performance differences being treated as ties. We consider language understanding (LU), commonsense reasoning (CR), question answering (QA), physics/biology/chemistry (PBC), math (Math), and medicine (Med) categories. Kendall’s $\tau$-b is averaged across all pairs of benchmarks that belong to two given categories. The diagonal represents the intra-category agreement and the others represent the inter-category agreement. train-before-test improves both intra- and inter-category ranking agreement in all instances.
Figure 14: Cross benchmark ranking agreement under direct evaluation, measured by Kendall’s $tau$-b with insignificant model comparisons treated as ties.
Figure 15: Cross benchmark ranking agreement under train-before-test, measured by Kendall’s $tau$-b with insignificant model comparisons treated as ties.
Appendix D Broader Impacts and Limitations
Due to resource constraints, we do not conduct exhaustive hyperparameter searches for each model and benchmark. Instead, we use the same set of hyperparameters for all models and benchmarks. We do not tune these hyperparameters for any given model or benchmark. Therefore, all models received the same amount of hyperparameter search budget, albeit small.
We restrict our analysis to benchmarks that have a train set. However, recent benchmarks seldom include a train split. Thus, most of the benchmarks considered in our analysis, while typically highly cited and influential, were released before 2022.
We do not anticipate any direct societal impacts from this work, such as potential malicious or unintended uses, nor do we foresee any significant concerns involving fairness, privacy, or security considerations. Additionally, we have not identified potential harms resulting from the application of this technology.

### v2

@@
 This dilemma of conflicting rankings hampers model selection, clouds model comparisons, and adds confusion to a growing ecosystem of competing models.
-Recent work attributed ranking disagreement to the phenomenon of training on the test task: As released, different models exhibit a different level of preparation for any given test task.
-A candidate solution to the problem is train-before-test: Give each model the same benchmark-specific finetuning before evaluation.
-Our primary contribution is a broad empirical evaluation of train-before-test across 24 benchmarks and 61 models.
-We show that train-before-test significantly improves ranking agreement consistently across all benchmarks.
-Whereas rankings have little external validity to start with, they enjoy a significant degree of external validity when applying train-before-test: Model rankings transfer gracefully from one benchmark to the other.
-Even within the same model family, train-before-test reduces strong ranking disagreement to near-perfect agreement.
-In addition, train-before-test reduces the model-score matrix to essentially rank one, revealing new insights into the latent factors of benchmark performance.
+In this paper, we take a different perspective on model comparison: instead of relying on out-of-the-box performance via direct evaluation, we compare model potential by providing each model with identical benchmark-specific fine-tuning before evaluation.
+We call this approach train-before-test.
+Our primary contribution is a comprehensive empirical evaluation of model potential across 24 benchmarks and 61 models.
+First, we demonstrate that model potential rankings obtained through train-before-test exhibit remarkable consistency across all benchmarks.
+Whereas traditional rankings demonstrate little external validity under direct evaluation, they enjoy a significant degree of external validity when applying train-before-test: model potential rankings transfer gracefully from one benchmark to another.
+Second, train-before-test restores the connection between perplexity and downstream task performance, lost under direct evaluation.
+Remarkably, even pre-finetuning perplexity of a base model predicts post-finetuning downstream performance, suggesting that ranking consistency reflects inherent model potential rather than fine-tuning artifacts.
+Finally, train-before-test reduces the model-score matrix to essentially rank one, indicating that model potential is dominated by one latent factor, uncovered by train-before-test.
 Our work supports the recommendation to make train-before-test a default component of LLM benchmarking [^2] Code is available at https://github.com/socialfoundations/lm-harmony..
 1 Introduction
-A single large language model can reliably solve many tasks.
-This impressive capability, however, has led to a crisis in benchmarking.
-A rapidly growing flurry of benchmarks tries to capture the numerous different downstream tasks that we might use a language model for.
-Unfortunately, different benchmarks generally give contradictory model rankings, making it difficult to compare and choose models.
-Traditionally, benchmarking compared models trained for the same task on the same training data.
-In contrast, most large language models train on proprietary data mixes that differ from one model to the other.
-Recent work showed that this leads to the problem of training on the test task: the extent to which a model has encountered data similar to the test task during training confounds model comparisons, rankings, and scaling laws kaplan2020scaling .
-Put simply, an otherwise worse model may have simply prepared better for a specific task.
-A candidate solution to the problem is to fine-tune each model on the same task-specific training data before evaluation.
-We call this form of evaluation train-before-test.
-We draw a distinction between direct evaluation and train-before-test.
-In direct evaluation, we evaluate the model as it ships out of the box.
-Under train-before-test, we fine-tune the model before evaluation.
-The point of fine-tuning is not to make the model better, but to level the playing field for evaluation.
-The hope is that train-before-test might level out task-specific model differences so as to reveal more robust model comparisons.
-Our primary contribution is a broad evaluation of train-before-test, which confirms that this is indeed the case.
+Existing language model benchmarks provide contradictory model rankings, even for benchmarks that aim to capture similar skills (liang2022holistic, ; open-llm-leaderboard, ; open-llm-leaderboard-v2, ).
+This inconsistency poses a serious challenge: how can we reliably compare, rank, and select models when different benchmarks yield conflicting information?
+While this ranking disagreement is often attributed to the diverse capabilities of large language models (Ruan2024ObservationalSL, ), it creates a conundrum in practice that muddles model development decisions (zhang2024inherent, ).
+Current evaluation methodology works from direct evaluation, probing models via black-box function calls.
+However, large language models are trained on diverse, often proprietary data mixes that vary significantly across models (llama, ; gemma, ; Guha2025OpenThoughtsDR, ).
+Recent work showed that this leads to the problem of training on the test task (dominguez2024training, ): the extent to which a model has encountered data similar to the test task during training confounds model comparisons, rankings, and scaling laws kaplan2020scaling .
+Put simply, an otherwise inferior model may have simply prepared better for a specific task.
+In this paper, we take a fresh perspective on evaluation methodology: in contrast with direct evaluation, we compare model potential by giving each model the same task-specific fine-tuning.
+We call this approach train-before-test.
+Its goal is to achieve valid model comparisons by ensuring that all models receive equal preparation for the test.
+We envision train-before-test as a tool for regret-free model selection for downstream applications.
+Increasingly, practitioners select one from many available models with the goal of adapting for a specific task.
+Under direct evaluation the best model to begin with may no longer be the best model after task-specific preparation.
+In contrast, we show that train-before-task yields model comparisons and rankings that enjoy broad external validity.
 1.1 Our Contributions
@@
 See Appendix C.1 for more details.
-In more detail, we make the following key contributions.
 Direct evaluation leads to ranking disagreement even between related tasks.
@@
 Moreover, rankings still strongly disagree when evaluating models from the same family.
-The situation presents a serious conundrum for model selection: Under direct evaluation, benchmarks fail to give reliable and actionable insights about which model to choose from a set of available alternatives.
-Train-before-test leads to ranking agreement.
-We broadly evaluate train-before-test on 24 benchmark datasets and 61 large language models.
-Train-before-test will fine-tune each model on the same task-relevant data before evaluation.
-Our results demonstrate that train-before-test drastically increases model ranking agreement across the board: Ranking agreement between benchmarks, measured by Kendall’s tau, improves for $95\%$ of the $\binom{24}{2}=276$ benchmark pairs.
-The average value of Kendall’s $\tau$ improves from $0.51$ to $0.74$.
+The situation presents a serious conundrum for model selection: Under direct evaluation, benchmarks fail to give reliable and actionable insights for model choosing among multiple alternatives.
+Train-before-test leads to consistent model potential rankings.
+We comprehensively evaluate train-before-test across 24 benchmark datasets and 61 large language models.
+By fine-tuning each model on identical task-relevant data before evaluation, we uncover remarkably consistent model potential rankings.
+Ranking agreement between benchmarks, measured by Kendall’s tau, improves for 274 out of 276 benchmark pairs, with the average Kendall’s $\tau$ increasing from 0.52 to 0.76.
 Figure 1 illustrates the result for one typical pair of benchmarks.
-Train-before-test aligns perplexity rankings with downstream tasks.
-Perplexity benchmarks used to be popular, but fell out of fashion because of the apparent disconnect between perplexity and downstream task performance (weiemergent, ; ganguli2022predictability, ; liu2023same, ).
+This consistency suggests that model potential, unlike out-of-the-box performance, has external validity (Salaudeen2025MeasurementTM, ) and transfers gracefully across different tasks.
+Model potential aligns perplexity rankings with downstream tasks.
+Perplexity benchmarks used to be popular, but fell out of fashion because of the apparent disconnect between perplexity and downstream task performance (weiemergent, ; ganguli2022predictability, ; liu2023same, ; Magnusson2023PalomaAB, ; Lourie2025ScalingLA, ).
 We indeed validate this disconnect when comparing model families under direct evaluation.
-In contrast, we demonstrate that under train-before-test evaluation, perplexity rankings align well with downstream task rankings.
-This creates an appealing consistency between model training objectives and downstream performance.
-Train-before-test sheds light on the latent factors of benchmark performance.
-Consider the large benchmark-model matrix, where each entry $(i,j)$ corresponds to the performance of model $j$ on a benchmark $i$.
+However, train-before-test restores this fundamental relationship in two ways.
+First, we show that post-fine-tuning perplexity rankings align well with post-fine-tuning downstream task rankings, creating consistency between training objectives and task performance.
+Second, and more remarkably, for base (non-instruction-tuned) models, even pre-fine-tuning perplexity predicts post-fine-tuning downstream performance.
+This suggests that the ranking consistency we observe reflects inherent model potential rather than artifacts of fine-tuning.
+Train-before-test sheds light on the latent factors of benchmark scores.
+Consider the large benchmark-model score matrix, where each entry $(i,j)$ corresponds to the performance of model $j$ on a benchmark $i$.
 Several works have considered this matrix for different reasons and found that it is approximately low rank (Ruan2024ObservationalSL, ; Owen2024HowPI, ; Burnell2023RevealingTS, ), but not quite.
 The first singular value is dominant and correlates with pre-training compute.
-However, the other components aren’t negligible and it wasn’t clear how to interpret them.
+However, the other components aren’t negligible, and their interpretation remains unclear.
 We show that train-before-test clarifies this state of affairs.
-After train-before-test, the benchmark-model matrix is essentially rank $1$.
-The first principal component accounts for $85\%$ of the explained variance across all models, and for $93\%$ of the variance for a single model family.
-This leads us to speculate that the lower components of the direct evaluation matrix correspond to how much a model was trained on a given kind of task.
-This would be a compelling reason why the lower singular values shrink under train-before-test.
+After train-before-test, the benchmark-model matrix is essentially rank one.
+The first principal component accounts for $86\%$ of the explained variance across all models, and for $93\%$ of the variance for a single model family.
+This suggests that model potential is dominated by a single latent factor, while the additional components observed in direct evaluation may reflect task-specific training exposure.
 2 Related Work
 Benchmarking has played a central role in the advancement of machine learning (liberman2010obituary, ; hardt2022patterns, ).
-While absolute model performance is often fragile to even seemingly minor changes in evaluation data (candela2009dataset, ; torralba2011unbiased, ; albadawy2018deep, ; taori2020measuring, ; Tsipras2020FromIT, ; miller2020effect, ), relative model performance—that is, model rankings—tends to transfer surprisingly well across classical benchmarks (Yadav2019ColdCT, ; recht2019imagenetclassifiersgeneralizeimagenet, ; miller2020effect, ).
+While absolute model performance is often fragile to even seemingly minor changes in evaluation data (candela2009dataset, ; torralba2011unbiased, ; albadawy2018deep, ; taori2020measuring, ; Tsipras2020FromIT, ), relative model performance—that is, model rankings—tends to transfer surprisingly well across classical benchmarks (Yadav2019ColdCT, ; recht2019imagenetclassifiersgeneralizeimagenet, ; miller2020effect, ).
 For instance, prior work (kornblith2019better, ; barbu2019objectnet, ) has shown that model rankings on ImageNet (deng2009imagenet, ) also transfer to other image classification and object recognition benchmarks.
-Moreover, Salaudeen and Hardt (2024, (salaudeen2024imagenot, )) demonstrated that ImageNet rankings remain robust even under major dataset variations (salaudeen2024imagenot, ).
+Moreover, Salaudeen and Hardt (2024, (salaudeen2024imagenot, )) demonstrated that ImageNet rankings remain robust even under major dataset variations.
 This transferability of model rankings is highly desirable, as it indicates that progress on specific benchmarks reliably reflects broader scientific advancements (liao2021we, ; hardt2025emerging, ).
 However, the emergence of foundation models has dramatically transformed the benchmarking landscape compared to the ImageNet era (liang2022holistic, ; srivastava2022beyond, ; Weidinger2025TowardAE, ).
-With huge training costs and much improved capabilities (qwen, ; gemma, ; llama, ; Ramesh2021ZeroShotTG, ; Anil2023GeminiAF, ; openai2023GPT4, ), practitioners now lean towards directly evaluating LLMs across a wide range of different benchmarks, in the hope of obtaining a more comprehensive assessment of their capabilities (liang2022holistic, ; bbh, ; mmlu, ; open-llm-leaderboard, ; open-llm-leaderboard-v2, ).
-This shift introduces new challenges, as model rankings across different tasks may vary significantly (liang2022holistic, ; huan2025does, ; lourie2025scaling, ).
-Zhang and Hardt (2024, (zhang2024inherent, )) draw an analogy between multi-task benchmarks and voting systems (arrow2012social, ), revealing that a multi-benchmarking approach with diverse rankings inherently lacks robustness to minor changes and thus cannot provide a stable unified ranking.
+With huge training costs and much improved capabilities (qwen, ; llama, ; Ramesh2021ZeroShotTG, ; Anil2023GeminiAF, ; openai2023GPT4, ), practitioners now lean towards directly evaluating LLMs across a wide range of different benchmarks, in the hope of obtaining a more comprehensive assessment of their capabilities (liang2022holistic, ; bbh, ; mmlu, ; open-llm-leaderboard, ; open-llm-leaderboard-v2, ).
+This shift introduces new challenges, as model rankings across different tasks may vary significantly (huan2025does, ; lourie2025scaling, ).
+Zhang and Hardt (2024, (zhang2024inherent, )) draw an analogy between multi-task benchmarks and voting systems (arrow2012social, ), revealing that a multi-task benchmarking approach with diverse rankings inherently lacks robustness to minor changes and thus cannot provide a stable unified ranking.
 This lack of unified ranking is sometimes seen as a desirable feature within the community (liang2022holistic, ).
@@
 Given the resource-intensive nature of LLM development (guo2025deepseek, ), it is impractical to release a different model for every potential use case.
-Moreover, developers typically aim to create a general-purpose model (qwen, ; llama, ; gemma, ); however, such a desideratum is often difficult to reliably measure due to the inconsistent rankings observed across benchmarks.
+Moreover, developers typically aim to create a general-purpose model (qwen, ; llama, ); however, such a desideratum is often difficult to reliably measure due to the inconsistent rankings observed across benchmarks.
 Second, we demonstrate in this paper that benchmarks within the same task category can still exhibit substantial discrepancies in model rankings.
@@
 (2024, (dominguez2024training, )) show that models vary in their degree of preparedness for popular benchmarks.
-Building on this idea, we introduce the notion of train-before-test, wherein we fine-tune each model on the corresponding training set to get every model well-prepared.
-We then investigate how this approach improves the consistency of rankings across benchmarks and discuss its implications for future benchmarking practices.
+Building on this idea, we introduce the notion of train-before-test, wherein we fine-tune each model on the corresponding training set so every model arrives well prepared.
+While extensive literature exists on investigating different fine-tuning strategies for LLMs (Lin2024SelectingLL, ; Zeng2025LENSLLMUF, ; Lester2021ThePO, ), this lies outside the scope of our investigation.
+Instead, we apply standardized fine-tuning (peft, ) as an evaluation tool to give all models equivalent preparation before testing.
+Rather than studying models from the same family with varying pre-training compute (Zhang2024WhenSM, ; kaplan2020scaling, ), our experiments cover 61 models from six families and 24 tasks from different categories.
+We study how train-before-test improves ranking consistency across benchmarks and its implications for benchmarking practices.
 3 Experiments
 In this section, we examine the cross-benchmark ranking agreement of 61 language models across 24 benchmarks.
-We find that ranking agreement tends to be low, with an average Kendall’s $\tau$ of 0.51.
+We find that ranking agreement tends to be low, with an average Kendall’s $\tau$ of 0.52.
 We then examine benchmark agreement under a different benchmarking methodology, which we refer to as train-before-test.
 Specifically, we fine-tune on a benchmark’s train set prior to evaluating on said benchmark.
-Compared to direct evaluation, train-before-test improves cross-benchmark ranking agreement on 95% of the benchmark pairs considered.
-The improvements in ranking agreement are typically large, with the average Kendall’s $\tau$ increasing to 0.74.
+Compared to direct evaluation, train-before-test improves cross-benchmark ranking agreement on almost all benchmark pairs considered.
+The improvements in ranking agreement are typically large, with the average Kendall’s $\tau$ increasing to 0.76.
 We additionally find that train-before-test greatly improves the agreement between perplexity rankings and downstream benchmarks.
 This result holds consistently across three general domain corpus, newly collected from Wiki, arXiv, and StackExchange.
-The average Kendall’s $\tau$ improves from 0.47 to 0.73 with train-before-test, leading to much better consistency between the training objective and downstream benchmark performance.
+We retrained content only from 2025 to ensure models had not seen those texts during pretraining.
+The average Kendall’s $\tau$ between perplexity ranking and 24 downstream task rankings improves from 0.48 to 0.74 with train-before-test, leading to much better consistency between the training objective and downstream benchmark performance.
+Moreover, for base models, pre-fine-tuning perplexity ranking remains consistent with post-fine-tuning downstream rankings (average Kendall’s $\tau$ = 0.78).
+This consistency does not hold for instruction-tuned models.
 Finally, we discuss the implications of increased cross-benchmark agreement.
@@
 We then analyze its principal components both under direct model evaluation and train-before-test.
-We find that train-before-test greatly increases the share of variance explained by the first principal component (PC1), from 69% for direct evaluation to 85% for train-before-test.
+We find that train-before-test greatly increases the share of variance explained by the first principal component (PC1), from 70% for direct evaluation to 86% for train-before-test.
 In both cases, PC1 aligns well with model pre-training compute.
 By further controlling for model family and only considering Qwen models, we show that the explained variance ratio of PC1 further increases to 93%, making the model-score matrix essentially rank one.
+3.1 Experiment Setting
 Table 1: We categorize benchmarks into language understanding (LU), commonsense reasoning (CR), question answering (QA), physics/biology/chemistry (PBC), math (Math), and medicine (Med).
@@
 Med MedMCQA (medmcqa, ), HeadQA (headqa, )
-Figure 2: Mean ranking agreement between each benchmark and all others.
-We calculate Kendall’s $\tau$ between each benchmark and every other benchmark, and then average the results.
-Compared to direct evaluation, train-before-test consistently improves ranking agreement, often by a large margin.
-A detailed comparison of Kendall’s $\tau$ values for every benchmark pair is provided in Appendix B.1.
-On average, the overall average Kendall’s $\tau$ is 0.51 for direct evaluation and 0.74 for train-before-test.
-(a) Direct evaluation.
-(b) Train-before-test.
-Figure 3: Cross-category ranking agreement for direct evaluation (left) and train-before-test (right).
-We categorize benchmarks into language understanding (LU), commonsense reasoning (CR), question answering (QA), physics/biology/chemistry (PBC), math (Math), and medicine (Med), see Table 1.
-Kendall’s $\tau$ is averaged across all pairs of benchmarks that belong to two given categories.
-The diagonal entries represent intra-category agreement and the other entries represent inter-category agreement.
-Train-before-test improves both intra- and inter-category ranking agreement in all instances.
-3.1 Experiment Setting
 Benchmark selection.
@@
 Generation-based benchmarks are often computationally intensive to evaluate, as base models typically generate text until reaching their maximum sequence length.
-We select only NQ-Open and GSM8K from the generation-based benchmarks, excluding the remaining seven due to computational constraints.
+These benchmarks are also over-challenging for smaller models with limited parameters, such as GPT-2 (gpt2, ).
+Therefore, we select only NQ-Open and GSM8K from the generation-based benchmarks.
 Among the likelihood-based benchmarks, we further exclude six due to observed anomalies during fine-tuning, such as a lack of performance improvement in over 20% of models.
-See Appendix A.1 for additional details.
+See Appendix A.1 for details.
 Our final selection consists of 24 benchmarks covering diverse domains and task types.
@@
 To save computational resources, we cap the number of training data at 50,000, validation data at 1,000, and testing data at 10,000.
+Table 2: Models considered, categorized by model family.
+Family Model Name Suffix
+Llama-
+3-8B, 3.1-8B, 3.2-1B, 3.2-3B, 3-8B-IT, 3.1-8B-IT, 3.2-1B-IT, 3.2-3B-IT
+Qwen-
+1.5-0.5B, 1.5-1.8B, 1.5-4B, 1.5-7B, 1.5-14B, 2-0.5B, 2-1.5B, 2-7B, 2.5-0.5B, 2.5-1.5B, 2.5-3B, 2.5-7B,
+2.5-14B, 1.5-0.5B-IT, 1.5-1.8B-IT, 1.5-4B-IT, 1.5-7B-IT, 1.5-14B-IT, 2-0.5B-IT, 2-1.5B-IT, 2-7B-IT,
+2.5-0.5B-IT, 2.5-1.5B-IT, 2.5-3B-IT, 2.5-7B-IT, 2.5-14B-IT
+Gemma-
+2B, 7B, 2-2B, 2-9B, 2B-IT, 7B-IT, 2-2B-IT, 2-9B-IT
+GPT2-
+124M, 335M, 774M, 1.5B
+Pythia-
+70M, 160M, 410M, 1B, 1.4B, 2.8B, 6.9B, 12B
+Yi-
+6B, 9B, 6B-IT, 1.5-6B, 1.5-9B, 1.5-6B-IT, 1.5-9B-IT
 Model selection.
 We consider 61 language models across six model families: Llama (llama, ), Qwen (qwen, ), Gemma (gemma, ), Pythia (pythia, ), GPT-2 (gpt2, ) and Yi (yi, ).
-Due to computational constraints, we select models with fewer than 14B parameters.
+Due to computational constraints, we select models with no more than 14B parameters.
 See Table 2 for the full list.
@@
 For direct evaluation, we simply evaluate the model as it is.
-For train-before-test, we fine-tune models for five epochs, and select the best performing checkpoint based on performance on a separate validation set.
-We therefore fine-tune a total of $61$ models on $24$ benchmarks, totaling $61\times 24=1,464$ fine-tuned models.
-We use parameter-efficient fine-tuning (PEFT) (Hu2021LoRALA, ; peft, ), with fixed hyperparameters.
+For train-before-test, we fine-tune models for five epochs using learning rates in $\{1\mathrm{e}-5,2\mathrm{e}-5,5\mathrm{e}-5\}$, separately.
+The best performing checkpoint is then selected based on performance on a separate validation set, yielding $61\times 24=1,464$ fine-tuned models in total.
+We use parameter-efficient fine-tuning (PEFT) (Hu2021LoRALA, ; peft, ).
 See more details in Appendix A.2.
@@
 We then measure the ranking correlation across pairs of benchmarks using Kendall’s $\tau$ (kendall1938new, ).
+Figure 2: Mean ranking agreement between each benchmark and all others.
+We calculate Kendall’s $\tau$ between each benchmark and every other benchmark, and then average the results.
+Compared to direct evaluation, train-before-test consistently improves ranking agreement, often by a large margin.
+A detailed comparison of Kendall’s $\tau$ values for every benchmark pair is provided in Appendix B.1.
+On average, the overall average Kendall’s $\tau$ is 0.52 for direct evaluation and 0.76 for train-before-test.
+(a) Direct evaluation.
+(b) Train-before-test.
+Figure 3: Cross-category ranking agreement for direct evaluation (left) and train-before-test (right).
+We categorize benchmarks into language understanding (LU), commonsense reasoning (CR), question answering (QA), physics/biology/chemistry (PBC), math (Math), and medicine (Med), see Table 1.
+Kendall’s $\tau$ is averaged across all pairs of benchmarks that belong to two given categories.
+The diagonal entries represent intra-category agreement and the others represent inter-category agreement.
+Train-before-test improves both intra- and inter-category ranking agreement in all instances.
 3.2 Downstream Ranking Agreement
-As depicted in Figure 2, direct evaluation shows only modest ranking agreement between the 24 benchmarks, with an average Kendall’s $\tau$ ranking correlation of 0.51.
-This lack of agreement across benchmarks complicates model assessment and makes it challenging to aggregate results into a meaningful overall ranking.
+As depicted in Figure 2, direct evaluation shows only modest ranking agreement between the 24 benchmarks, with an average Kendall’s $\tau$ of 0.52.
+This lack of agreement across benchmarks complicates model assessment and makes it challenging to aggregate results into a meaningful overall ranking (zhang2024inherent, ).
 In contrast, the train-before-test methodology leads to a substantial improvement in ranking agreement.
-Under this approach, 95% of model pairs show higher Kendall’s $\tau$ scores, with the average $\tau$ rising from 0.51 to 0.74.
-This improvement in ranking consistency is highly notable, as it indicates that models that excel on any given benchmark are also likely to excel in many others, thereby simplifying the model selection process for developers.
-Notably, benchmarks that appeared to be outliers under direct evaluation, such as MRPC and WiC, demonstrate much greater ranking consistency under train-before-test.
-For example, the average Kendall’s $tau$ between MRPC and all other benchmarks improves dramatically, from 0.23 to 0.73.
+Under this approach, 274 out of 276 benchmark pairs show higher Kendall’s $\tau$ scores, with the average $\tau$ rising from 0.52 to 0.76.
+This stronger consistency suggests that model potential ranking on one benchmark is likely to generalize to others, including practitioners’ own cases, which simplifies model comparison and selection.
+Notably, benchmarks that appeared to be outliers under direct evaluation, such as NQ-Open and MRPC, demonstrate much greater ranking consistency under train-before-test.
+For example, the average Kendall’s $\tau$ between NQ-Open and all other benchmarks improves from 0.23 to 0.74.
 We further split all benchmarks into six categories (e.g., language understanding, math), see Table 1.
 For each category pair, we report in Figure 3 the intra-category average ranking correlations and inter-category average ranking correlations across all relevant benchmark pairs.
-Consistent with our previous findings, we observe reasonably poor ranking agreements across benchmark categories.
-While one might expect high intra-category agreement–after all, tasks within the same category tend to be relatively similar–direct evaluation results in low intra-category agreement in many cases.
-For example, the intra-category mean Kendall’s $\tau$ is 0.52 for language understanding and 0.56 for math.
+Consistent with our previous findings, we observe reasonably poor ranking agreements across categories under direct evaluation.
+While one might expect high intra-category agreement—after all, tasks within the same category tend to be relatively similar—direct evaluation results in low intra-category agreement in many cases.
+For example, the intra-category mean Kendall’s $\tau$ is 0.54 for language understanding and 0.55 for math.
 This further underscores the difficulty of selecting models based on direct evaluation.
-Even if the goal is to choose a model that excels not across all tasks but within a specific domain, the low intra-category agreement makes this decision challenging.
-In contrast, train-before-test markedly boosts both intra- and inter-category consistency.
-For example, the intra-category mean Kendall’s $\tau$ for language understanding raises from 0.52 to 0.71, as well as from 0.56 to 0.82 for the math category.
+Even if the goal is to choose a model that excels within a specific domain, the low intra-category agreement makes this decision challenging.
+In contrast, train-before-test boosts both intra- and inter-category consistency.
+For example, the intra-category mean Kendall’s $\tau$ for language understanding raises from 0.52 to 0.75, as well as from 0.55 to 0.84 for the math category.
 Moreover, agreement between categories is often nearly as high as agreement within categories.
-This suggests that models performing well in one domain tend to also perform well across other domains.
+This suggests that models with higher potential in one domain tend to also perform well across other domains after adaptation.
 3.3 Perplexity Agreement
-We now study the agreement between benchmark rankings and rankings based on perplexity on general domain corpora.
+We now study the agreement between downstream benchmark rankings and perplexity rankings on general domain corpora.
 To do so, we collect three corpora from Wikipedia, StackExchange, and arXiv, retaining only contents from 2025.
-More specifically, we collect 3,366 documents for Wiki, 6,001 for StackExchange and 44,384 documents for arXiv.
+Because all models used were released before 2025, they could not have seen these texts during pretraining.
+Specifically, we collect 3,366 documents for Wiki, 6,001 for StackExchange and 44,384 for arXiv.
 These datasets are split into training, validation, and testing sets, in an 8:1:1 ratio.
@@
 Perplexity rankings are consistent with each other under both evaluation schemes, with an average Kendall’s $\tau$ of 0.76 and 0.78, respectively.
-However, for direct evaluation, agreement between perplexity rankings and downstream rankings is low, with an average Kendall’s $\tau$ of just 0.47.
-Fortunately, train-before-test results in much higher agreement between perplexity and downstream evaluations, increasing average Kendall’s $\tau$ to 0.73.
-Figure 5: Ranking agreement between perplexity rankings before fine-tuning and downstream benchmark rankings after fine-tuning, for base models (top) and instruction-tuned models (bottom).
-For base models, perplexity rankings before fine-tuning are consistent with downstream performance rankings after fine-tuning, with an average of Kendall’s $\tau$ 0.77.
-The same does not hold for instruction-tuned models, where the average Kendall’s $\tau$ between perplexity rankings and fine-tuning rankings is 0.5.
+However, for direct evaluation, agreement between perplexity rankings and downstream rankings is low, with an average Kendall’s $\tau$ of just 0.48.
+Fortunately, train-before-test results in much higher agreement between perplexity and downstream evaluations, increasing average Kendall’s $\tau$ to 0.74.
+Figure 5: Ranking agreement between perplexity rankings before fine-tuning (direct evaluation) and downstream benchmark rankings after fine-tuning (train-before-test) for base models (top) and instruction-tuned models (bottom).
+Unlike Figure 4 where both rankings in each comparison use the same evaluation scheme, here we test whether pre-fine-tuning perplexity can predict post-fine-tuning downstream performance.
+Base models show strong correlation (average Kendall’s $\tau$ = 0.78), suggesting perplexity is a good predictor of model potential.
+Instruction-tuned models show much weaker correlation (average Kendall’s $\tau$ = 0.51).
 We measure perplexity in bits per byte, using the lm-eval-harness library.
-We then construct models rankings based on the perplexity evaluations, and contrast the rankings with those of the downstream benchmarks considered in earlier sections.
+We then compute models rankings based on the perplexity evaluations, and compare the rankings with those of the downstream benchmarks considered in earlier sections.
 We exclude the four Gemma models from these results, as lm-eval-harness provides unreliable perplexity measurements for Gemma models due to its rolling window implementation.
-See Appendix B.2 for more details.
+See Appendix B.2 for details.
 The results are presented in Figure 4.
-In contrast to downstream tasks, perplexity rankings demonstrate strong agreement both for direct evaluation and train-before-test.
+In contrast to downstream tasks, perplexity rankings demonstrate strong agreement both under direct evaluation and train-before-test.
 Specifically, the average Kendall’s $\tau$ between the perplexity rankings is 0.76 for direct evaluation and 0.78 for train-before-test.
 We hypothesize that this reasonably strong agreement arises due to the smooth relationship between perplexity evaluations (brandfonbrener2024loss, ; mayilvahananllms, ).
-When comparing ranking agreement between perplexity evaluations and downstream benchmarks, we find that agreement is low under direct evaluation, with a mean Kendall’s $\tau$ of 0.47.
+When comparing ranking agreement between perplexity evaluations and downstream benchmarks, we find that agreement is low under direct evaluation, with a mean Kendall’s $\tau$ of 0.48.
 This lack of agreement is concerning, as it signals a disconnect between the language modelling pre-training objective and downstream benchmark performance.
-Fortunately, we find that our train-before-test methodology improves ranking agreement substantially, with the mean Kendall’s $\tau$ ranking correlation between perplexity rankings and benchmark rankings rising to 0.73.
+Fortunately, we find that our train-before-test methodology improves ranking agreement substantially, with the mean Kendall’s $\tau$ ranking correlation between perplexity rankings and benchmark rankings rising to 0.74.
 This finding is reassuring: a light amount of fine-tuning on task data is sufficient to align the language modeling training objective with downstream performance.
@@
 This suggests that, despite perplexity typically not being used for benchmarking purposes, it can be as effective a ranking metric as benchmark evaluations.
-Drawing inspiration from prior work (liu2023same, ; xia2023training, ; Gadre2024LanguageMS, ; benchmarkloss, ), we further examine the correlation between model rankings according to average perplexity across the three text corpora and average downstream performance across the 24 benchmarks.
-Gadre et al.
-(2024, (Gadre2024LanguageMS, )) show that, when models are trained on the same pretraining data, perplexity is well-correlated with aggregate benchmark performance.
+Drawing inspiration from prior work (liu2023same, ; xia2023training, ; Gadre2024LanguageMS, ; benchmarkloss, ; Zhang2024WhenSM, ), we further examine the correlation between model rankings according to average perplexity across the three text corpora and average downstream performance across the 24 benchmarks.
+Zhang and Hardt (2024, (Gadre2024LanguageMS, )) show that, when models are trained on the same pretraining data, perplexity is well-correlated with aggregate benchmark performance.
 Our setup differs in that we consider a diverse set of model families, each trained on different pretraining data.
-Under direct evaluation, we find that the ranking correlation is modest, with a Kendall’s $\tau$ of only 0.56.
+Under direct evaluation, we find that the ranking correlation is modest, with a Kendall’s $\tau$ of only 0.55.
 We hypothesize that this relatively weak agreement is due to differences in pretraining data and instruction tuning, resulting in varying levels of exposure to benchmark tasks during training (dominguez2024training, ).
-Fortunately, when applying our train-before-test methodology, the ranking correlation between average perplexity and average downstream performance improves substantially, with Kendall’s $\tau$ increasing from 0.56 to 0.82.
+Fortunately, when applying our train-before-test methodology, the ranking correlation between average perplexity and average downstream performance improves substantially, with Kendall’s $\tau$ increasing from 0.55 to 0.84.
 We additionally examine the agreement between perplexity prior to fine-tuning and downstream task performance after fine-tuning.
@@
 We plot such ranking agreement in Figure 5, dividing models into base models and instruction-tuned models.
-We find that, for base models, perplexity prior to fine-tuning is a strong indicator of task performance after fine-tuning.
+For base models, perplexity prior to fine-tuning is a strong indicator of model potential on downstream tasks, with an average Kendall’s $\tau$ of 0.78.
 This indicates that, for base models, direct evaluation of perplexity is already a reasonably reliable metric for ranking models.
-However, the same does not hold for instruction-tuned models.
-We find that instruction-tuning renders perplexity rankings unreliable, as ranking agreement is low across the board.
-This is to be expected: instruction fine-tuning tends to increase both benchmark performance and perplexity on general text corpora, thus clouding the relationship between perplexity and downstream evaluations.
-Fortunately, as shown earlier, train-before-test restores high ranking agreement between perplexity evaluations and downstream task performance.
+Moreover, it indicates that the ranking consistency we observe reflects inherent model potential rather than artifacts introduced by fine-tuning.
+However, the same does not hold for instruction-tuned models (average Kendall’s $\tau$ = 0.51).
+Instruction-tuning renders perplexity rankings unreliable, as ranking agreement is low across the board.
+This is to be expected: instruction fine-tuning tends to increase both benchmark performance ($\uparrow$) and perplexity ($\downarrow$) on general text corpora, thus clouding the relationship between perplexity and downstream evaluations.
+Fortunately, as shown earlier, train-before-test restores high ranking agreement between perplexity evaluations and downstream performance.
 Figure 6: Explained variance ratios of the top five principal components of the benchmark score matrix, under direct evaluation (left) and train-before-test (right).
-Train-before-test substantially increases the amount of variance explained by the first principal component, from 69% to 85%
+Train-before-test substantially increases the amount of variance explained by the first principal component, from 70% to 86%
 3.4 Low-Ranked Model Score Matrix
-So far, we have shown that the train-before-test methodology significantly improves cross-benchmark ranking agreement.
+So far, we have shown that comparing model potential using the train-before-test yields consistent rankings across benchmarks.
 We now examine the implications of this finding by analyzing the resulting matrix of model scores, where each entry $(i,j)$ corresponds to the performance of model $j$ on a benchmark $i$.
-We use Principal Component Analysis (PCA) to examine the structure of the matrix of model scores.
+We use Principal Component Analysis (PCA) to examine the structure of the model score matrix.
+Figure 7: PC1 scores under train-before-test align with the pre-training compute.
 Figure 6 shows the explained variance ratios of the first five principal components.
@@
 A similar trend is observed for train-before-test scores, where the first five components explain 97% of the variance.
-Notably, under train-before-test, the first principal component (PC1) captures a significantly larger share of the variance: 85%, compared to 69% for direct evaluation.
-Prior works interpret PC1 scores as an indication of general capability, with later principal components denoting domain-specific capabilities not captured by PC1 (e.g., reasoning ability, coding ability) (Ruan2024ObservationalSL, ; Burnell2023RevealingTS, ).
-Contrary to the popular notion that the performance of LLMs is governed by multiple axes of capability (Ruan2024ObservationalSL, ; Burnell2023RevealingTS, ), our results highlight the dominance of a single principal axis of general capability.
-Figure 7: PC1 scores under train-before-test align with the pre-training compute.
-It is of no surprise that PC1 positively correlates with pre-training compute, as shown in Figure 7222We only plot models whose number of training tokens is publicly available.
-See Table 4 for details., which have been identified as crucial to model performances (kaplan2020scaling, ).
-However, the exact interpretation of PC1 remains open to debate.
-It could be model scale (kaplan2020scaling, ), general capability (Ruan2024ObservationalSL, ), or another factor entirely.
-Nonetheless, it is evident that PC1 is strongly linked with model performance–particularly under the train-before-test methodology– and thus might serve as an effective target for guiding further model development.
+Notably, under train-before-test, the first principal component (PC1) captures a much larger share of the variance: 86%, compared to 70% for direct evaluation.
+Prior works interpret PC1 scores under direct evaluation as an indication of general capability, with later principal components denoting domain-specific capabilities not captured by PC1 (e.g., reasoning ability, coding ability) (Ruan2024ObservationalSL, ; Burnell2023RevealingTS, ).
+Unlike out-of-the-box performance, which is controlled by multiple factors (Ruan2024ObservationalSL, ; Burnell2023RevealingTS, ), model potential is dominated by one single principal axis.
+It is of no surprise that PC1 also positively correlates with pre-training compute, as shown in Figure 7222We only plot models whose number of training tokens is publicly available.
+See Table 4 for details., which have been identified as crucial to model performances (kaplan2020scaling, ; Hoffmann2022TrainingCL, ).
 See detailed PC1 scores in Figure 11.
@@
 We repeat our PCA analysis on the score matrix containing only Qwen models, depicted in Figure 8.
-Remarkably, we find that PC1 for train-before-test explains 93% of the variance, roughly as much as the variance explained by the top four principal components under direct evaluation.
+Remarkably, we find that PC1 for train-before-test explains 93% of the variance, roughly as much as the variance explained by the top five principal components under direct evaluation.
 That is, whereas for direct evaluation the score matrix is low-rank; train-before-test renders the score matrix essentially rank one.
 4 Discussion, limitations, and conclusion
-Train-before-test harmonizes model rankings across numerous benchmarks we evaluated.
-Put differently, it endows model rankings with external validity they lack under direct evaluation.
-In turn, this greatly increases the utility of benchmarks for reliable model comparisons and model selection.
-To play devil’s advocate, however, one might ask why do we need ranking consistency if we can just look at the benchmark that best fits a downstream application?
-First, we show that even benchmarks that aim to capture the same task give contradictory rankings.
-Second, there will never be a perfect match between a benchmark and the downstream application.
-We must therefore always rely on some form of external validity, specifically, that the ranking transfers gracefully to other settings.
-Third, practitioners typically specialize any chosen model to their application.
-This makes direct evaluation numbers irrelevant, anyway.
+Train-before-test fundamentally reframes how we interpret model evaluation.
+Whereas direct evaluation yields benchmark-specific rankings that often contradict one another, train-before-test harmonizes rankings across a wide array of tasks and datasets.
+This shift from measuring out-of-the-box performance to comparing achievable potential equips the community with a more stable and externally valid evaluation methodology.
+This emphasis on model potential is particularly valuable for scenarios involving model development and adaptation.
+Practitioners frequently need to make decisions during model development—selecting checkpoints mid–pre-training or choosing a base model for further instruction tuning or domain-specific adaptation.
+In these scenarios, direct evaluation, while useful for assessing deployment readiness, is of limited relevance and utility.
+A model that performs poorly on direct evaluation might excel when adapted to new tasks.
+Train-before-test, by contrast, shows that rankings on any task will also generalize to others, offering more promising guidance for model selection.
+One might argue that ranking consistency is unnecessary if we can simply choose benchmarks close to a given downstream application.
+However, our findings highlight three challenges with that view.
+First, even benchmarks that purport to measure the same skill (e.g., question answering) produce contradictory rankings under direct evaluation.
+Second, no benchmark perfectly captures the specifics of an application, making some degree of generalization unavoidable.
+Third, in real deployments, models are often adapted to varying degrees, making the potential the relevant signal for comparison.
 Limitations.
-Train-before-test requires that we fine-tune models on agreed upon task-specific data prior to evaluation.
-This certainly increases the cost of evaluation and might be too costly in some cases.
-On the other hand, it also points to interesting avenues for future work on how to implement train-before-test most efficiently.
+Train-before-test requires fine-tuning models on task-specific data before evaluation, which certainly increases the evaluation cost.
+However, this investment yields dividends through improved reliability.
+Our findings suggest that fewer benchmarks suffice under train-before-test, as rankings from one benchmark reliably transfer to others.
+This reduction in required evaluations can offset the per-benchmark cost increase.
 A second problem is that, unfortunately, many benchmarks no longer come with training data, making it more difficult to apply train-before-test.
@@
 Conclusion.
-On the whole, we believe that train-before-test is a sound recommendation for the evaluation ecosystem that can help address some of its serious problems.
+Overall, train-before-test complements existing evaluation practices by distinguishing between performance and potential.
+cImportantly, potential comparison is not intended to replace direct evaluation—both serve distinct purposes.
+Direct evaluation remains useful for understanding immediate deployment readiness, while potential comparison provides insights into adaptability and development prospects.
+Together, they offer a more complete picture of model capabilities.
+We believe that adopting train-before-test as a standard alongside direct evaluation can significantly improve the reliability, interpretability, and practical utility of the model evaluation ecosystem.
 Acknowledgement
@@
 A.1 Benchmark Selection
-Language model evaluations can be broadly classified into likelihood-based and generation-based.
+We begin our study with the lm-eval-harness package [26], which offers a comprehensive suite of language model benchmarks.
+To accommodate the train-before-test methodology which requires a dedicated training set for fine-tuning, we first identify benchmarks that provide at least 1,000 training examples.
+This yields a total of 37 benchmarks.
+These benchmarks can be broadly categorized as 28 likelihood-based and 9 generation-based benchmarks.
 Likelihood-based evaluations test for the likelihood of different completions given some input string; for example, different answer choices given a multiple-choice input question.
@@
 This is particularly true for base models, which are usually not trained for instruction following and therefore continue to generate tokens until hitting their maximum token limit.
-Due to computational constraints, we exclude seven generation-based benchmarks, Drop, CoQa, ReCoRD, bAbi, WebQA, TriviaQA and Fld-Default.
+These generation-based benchmarks are also over-challenging for smaller models with limited parameters, such as GPT-2 [64].
+Therefore, we exclude seven generation-based benchmarks, Drop, CoQa, ReCoRD, bAbi, WebQA, TriviaQA and Fld-Default.
 Nevertheless, we retain two widely used generation-based benchmarks, GSM8K and NQ-Open, in our experiments.
 We additionally excluded five benchmarks due to anomalies observed during fine-tuning: MedQA-4Options, LogiQA, Mutual, Mela-EN, and Swag.
-For these benchmarks, more than 20% of models showed no performance improvement after fine-tuning.
+For these benchmarks, more than 20% of models, most of which are small models with fewer than 0.5B parameters, showed no performance improvement after fine-tuning.
 We also excluded Paws-EN, as its corresponding model ranking under direct evaluation was negatively correlated (Kendall’s $\tau$ less than zero) with 23 out of 24 other benchmarks.
 We attribute this anomaly to the unusual prompting template used by lm-eval-harness.
-A.2 Model Selection
-See Table 2 for the complete list of models used in our experiments.
-Table 2: Models considered, categorized by model family.
-Family Model Name Suffix
-Llama- [29] 3-8B, 3.1-8B, 3.2-1B, 3.2-3B, 3-8B-IT, 3.1-8B-IT, 3.2-1B-IT, 3.2-3B-IT
-Qwen- [56] 1.5-0.5B, 1.5-1.8B, 1.5-4B, 1.5-7B, 1.5-14B, 2-0.5B, 2-1.5B, 2-7B, 2.5-0.5B, 2.5-1.5B, 2.5-3B, 2.5-7B, 2.5-14B, 1.5-0.5B-IT, 1.5-1.8B-IT, 1.5-4B-IT, 1.5-7B-IT, 1.5-14B-IT, 2-0.5B-IT, 2-1.5B-IT, 2-7B-IT, 2.5-0.5B-IT, 2.5-1.5B-IT, 2.5-3B-IT, 2.5-7B-IT, 2.5-14B-IT
-Gemma- [71] 2B, 7B, 2-2B, 2-9B, 2B-IT, 7B-IT, 2-2B-IT, 2-9B-IT
-GPT2- [57] 124M, 335M, 774M, 1.5B
-Pythia- [8] 70M, 160M, 410M, 1B, 1.4B, 2.8B, 6.9B, 12B
-Yi- [82] 6B, 9B, 6B-IT, 1.5-6B, 1.5-9B, 1.5-6B-IT, 1.5-9B-IT
-A.3 Evaluation Setup
+Our final selection includes 24 benchmarks: MNLI [85], QNLI [65], RTE [17, 30, 7], CoLA [81], SST-2 [73], MRPC [19], QQP, WiC [63], ANLI [59], Winogrande [46], CommonsenseQA [76], Hellaswag [90], Social-IQA [71], OpenBookQA [57], NQ-Open [44], BoolQ [14], ARC-Easy, ARC-Challenge [15], SciQ [84], PIQA [9], MathQA [3], GSM8K [16], MedMCQA [62], HeadQA [80].
+A.2 Evaluation Setup
 For our train-before-test evaluations, we fine-tune each model for five epochs and select the best-performing checkpoint based on evaluations on a separate validation set.
-To reduce memory consumption, we employ parameter-efficient fine-tuning (PEFT) [34, 47], allowing us to conduct most experiments on Quadro RTX 6000, Tesla V100-SXM2-32GB and NVIDIA A100-SXM4-80GB GPUs.
-We use fixed hyperparameters across all experiments.
-These hyperparameters were not tuned for any particular model or benchmark.
-Specifically, we use AdamW with a learning rate of 2e-5 and weight decay of 0.01.
-For PEFT, we use a LoRA configuration with rank 8, $\alpha=32$, and dropout 0.1.
-For models that do not exhibit performance improvements after fine-tuning, we simply report their pre-fine-tuning results.
-Since all training datasets considered in our work are publicly available, models may have already trained on them during pre-training.
-In certain cases, this may lead to further fine-tuning not being beneficial.
-For instruction-tuned models, we assess performance both with and without applying chat templates and use the best performing one under direct evaluation.
-In the train-before-test setting, we similarly fine-tune two variants: one where the training data are formatted with chat templates and one without.
-We similarly select the best performing approach.
-We observe that not using a chat template often leads to better performance than using a chat template, both for direct evaluation and train-before-test.
+We use the AdamW optimizer with a weight decay of 0.01.
+For each model-benchmark combination, we perform a hyperparameter search over three learning rates $\{1e-5,2e-5,5e-5\}$ and select the optimal one based on validation performance.
+To reduce memory consumption, we employ parameter-efficient fine-tuning (PEFT) [38, 55], We use a LoRA configuration with rank 8, $\alpha=32$, and dropout 0.1.
+Most of our experiments are conducted on Quadro RTX 6000, Tesla V100-SXM2-32GB and NVIDIA A100-SXM4-80GB GPUs.
+In cases where models show no performance improvement after fine-tuning, we report their pre-fine-tuning results.
+This scenario is rare and typically occur with smaller models (less than 500M parameters) that lack the capacity to perform certain tasks, resulting in near-random performance both before and after fine-tuning.
+Additionally, since all training datasets in our study are publicly available, some models may have encountered this data during pre-training, potentially limiting the benefits of additional fine-tuning.
+For instruction-tuned models, we evaluate performance both with and without chat templates, selecting the configuration that yields better results.
+Specifically, during direct evaluation, we assess model performance on the validation set under both conditions and apply the better-performing configuration to the test set.
+In the train-before-test setting, we similarly fine-tune two variants: one with training data formatted using chat templates and one without.
+We then select the approach that achieves the best performance on the validation set for final evaluation.
 Appendix B Additional Experiment Results
@@
 Consequently, the average document length is 163 words for arXiv, 250 words for StackExchange, and 1,502 words for Wikipedia.
-We exclude Gemma models from our perplexity agreement experiments, as lm-eval-harness provides unreliable perplexity measurements for Gemma models333 See discussion at https://github.com/huggingface/transformers/issues/29250.
+We exclude Gemma models from our perplexity agreement experiments, as lm-eval-harness provides unreliable perplexity measurements for Gemma models [^3] See discussion at https://github.com/huggingface/transformers/issues/29250.
 .
@@
 Figure 11: PC1 scores under train-before-test.
-[…]
+Table 4: The models used in Figure 7.
+The number of training tokens of these models is publicly available.
+We compute the number of pre-training FLOPs as $6\times\text{\#Parameters}\times\text{\#Tokens}$.
 Model #Parameters (B) #Tokens (B) #FLOPs (10^18)
@@
 We observe that accounting for statistical significance in models’ performance differences leads to slightly higher ranking correlations, as measured by Kendall’s $\tau$-b.
-For direct evaluation, average agreement increases from 0.51 to 0.57.
-For train-before-test, average agreement increases from 0.74 to 0.77.
-Therefore, train-before-test continues to lead to large improvements in raking agreement (from Kendall’s $\tau$-b 0.57 to 0.77).
+For direct evaluation, average agreement increases from 0.52 to 0.58.
+For train-before-test, average agreement increases from 0.76 to 0.77.
+Therefore, train-before-test continues to lead to large improvements in raking agreement (from Kendall’s $\tau$-b 0.58 to 0.77).
 Input: Model performance scores and standard errors
@@
 $(order_{1},order_{2})\leftarrow$ parallel_greedy_rank(models, $G_{1}$, $G_{2}$, score1, score2)
-rank1[$m$] $=$ position of $m$ in $order_{1}$ (starting from 1)
-rank2[$m$] $=$ position of $m$ in $order_{2}$ (starting from 1)
+rank1[$m$] == position of $m$ in $order_{1}$ (starting from 1)
+rank2[$m$] == position of $m$ in $order_{2}$ (starting from 1)
 return rank1, rank2
 Algorithm 3 rank_models(score1, stderr1, score2, stderr2)
-Figure 12: Mean ranking agreement between each benchmark and all others, measured by Kendall’s $tau$-b, with non-statistically significant performance differences being treated as ties.
+Figure 12: Mean ranking agreement between each benchmark and all others, measured by Kendall’s tautau-b, with non-statistically significant performance differences being treated as ties.
 We calculate Kendall’s $\tau$-b between each benchmark and every other one, and then average.
 Compared to direct evaluation, train-before-test consistently improves ranking agreement–often by a large margin.
-On average, the overall average Kendall’s $\tau$ is 0.57 for direct evaluation and 0.77 for train-before-test.
+On average, the overall average Kendall’s $\tau$ is 0.58 for direct evaluation and 0.77 for train-before-test.
 (a) Direct evaluation.
 (b) Train-before-test.
-Figure 13: Cross-category ranking agreement for direct evaluation (left) and train-before-test (right), measured by Kendall’s $tau$-b, with non-statistically significant performance differences being treated as ties.
+Figure 13: Cross-category ranking agreement for direct evaluation (left) and train-before-test (right), measured by Kendall’s tautau-b, with non-statistically significant performance differences being treated as ties.
 We consider language understanding (LU), commonsense reasoning (CR), question answering (QA), physics/biology/chemistry (PBC), math (Math), and medicine (Med) categories.
@@
 train-before-test improves both intra- and inter-category ranking agreement in all instances.
-Figure 14: Cross benchmark ranking agreement under direct evaluation, measured by Kendall’s $tau$-b with insignificant model comparisons treated as ties.
-Figure 15: Cross benchmark ranking agreement under train-before-test, measured by Kendall’s $tau$-b with insignificant model comparisons treated as ties.
-Appendix D Broader Impacts and Limitations
-Due to resource constraints, we do not conduct exhaustive hyperparameter searches for each model and benchmark.
-Instead, we use the same set of hyperparameters for all models and benchmarks.
-We do not tune these hyperparameters for any given model or benchmark.
-Therefore, all models received the same amount of hyperparameter search budget, albeit small.
-We restrict our analysis to benchmarks that have a train set.
-However, recent benchmarks seldom include a train split.
-Thus, most of the benchmarks considered in our analysis, while typically highly cited and influential, were released before 2022.
-We do not anticipate any direct societal impacts from this work, such as potential malicious or unintended uses, nor do we foresee any significant concerns involving fairness, privacy, or security considerations.
-Additionally, we have not identified potential harms resulting from the application of this technology.
+Figure 14: Cross benchmark ranking agreement under direct evaluation, measured by Kendall’s tautau-b with insignificant model comparisons treated as ties.
+Figure 15: Cross benchmark ranking agreement under train-before-test, measured by Kendall’s tautau-b with insignificant model comparisons treated as ties.